OpenStudy (anonymous):

can someone tell me if i got this right, -32y is the answer for -4y+28 ?

OpenStudy (anonymous):

are u trying to solve for y?

OpenStudy (anonymous):

trying to find the greatest common factor

OpenStudy (anonymous):

-4y+28=0 -4y=-28 divide both sides by -4 y=7

OpenStudy (anonymous):

o ok ty

OpenStudy (phi):

The greatest common factor of -4y+28 is 4 this is the biggest number (greatest number) that divides both (common to both) terms -4y/4 = -y (divides evenly) 28/4= 7 (divides evenly) notice that -4y/-32y = 1/8 (not a whole number) and 28/(-32y)= -7/(8y) not even close to a whole number.